### v1

# NSmark: Null Space Based Black-box Watermarking Defense Framework for Pre-trained Language Models

Abstract
Pre-trained language models (PLMs) have emerged as critical intellectual property (IP) assets that necessitate protection. Although various watermarking strategies have been proposed, they remain vulnerable to Linear Functionality Equivalence Attack (LFEA), which can invalidate most existing white-box watermarks without prior knowledge of the watermarking scheme or training data. This paper further analyzes and extends the attack scenarios of LFEA to the commonly employed black-box settings for PLMs by considering Last-Layer outputs (dubbed LL-LFEA). We discover that the null space of the output matrix remains invariant against LL-LFEA attacks. Based on this finding, we propose NSmark, a task-agnostic, black-box watermarking scheme capable of resisting LL-LFEA attacks. NSmark consists of three phases: (i) watermark generation using the digital signature of the owner, enhanced by spread spectrum modulation for increased robustness; (ii) watermark embedding through an output mapping extractor that preserves PLM performance while maximizing watermark capacity; (iii) watermark verification, assessed by extraction rate and null space conformity. Extensive experiments on both pre-training and downstream tasks confirm the effectiveness, reliability, fidelity, and robustness of our approach. Code is available at https://github.com/dongdongzhaoUP/NSmark.
NSmark: Null Space Based Black-box Watermarking Defense Framework for Pre-trained Language Models
Haodong Zhao1,2, Jinming Hu1, Peixuan Li1, Fangqi Li1, Jinrui Sha1, Peixuan Chen2, Zhuosheng Zhang1,∗, Gongshen Liu1,thanks: Corresponding authors. 1Shanghai Jiao Tong University 2Tencent 1 {zhaohaodong, hujinming, peixuan.li, solour_lfq, sjrgckym, zhangzs, lgshen}@sjtu.edu.cn 2 noahchen@tencent.com
1 Introduction
During past decades, with the rapid development of artificial intelligence (AI), pre-trained language models (PLMs) have achieved superior performance and been applied in a wide range of fields. At the same time, training high-performance PLMs requires a large amount of data and computing resources, thus PLMs can be regarded as valuable intellectual property (IP). With the deployment and application of machine learning as a service (MLaaS) platforms, companies sell well-trained PLMs as commodities. Once sold models are illegally stolen, distributed or resold, the rights of the model owner will be seriously infringed. Therefore, it is necessary to protect IP of PLMs.
Figure 1: Illustration of different watermark schemes against LFEA/LL-LFEA.
Watermarking techniques have been widely applied for protecting IP of deep learning models Chen et al. (2024); He et al. (2024). By incorporating identifiable information, these methods serve as a means to verify model ownership and provide proof of authenticity. Existing schemes can be divided into white-box and black-box categories according to whether the model parameters need to be accessed in verification. Among them, black-box schemes are more applicable when model parameters are inaccessible in most real-life scenarios, such as deployment as API.
However, protecting the IP of PLMs through watermarking presents significant challenges. Since PLMs can be deployed for various downstream tasks post-training Zhang et al. (2023), it is crucial for watermark schemes to be task-independent. Besides, recent studies have revealed vulnerabilities and shortcomings in existing watermarking techniques Li et al. (2023a). The proposed Linear Functionality Equivalence Attack (LFEA) is easy to conduct and can invalidate most existing white-box watermarks without knowledge of the watermarking scheme or the training data by exploiting linear invariance. Considering Last-Layer outputs, we further analyze and expand the attack scenarios black-box settings using model outputs (dubbed LL-LFEA).
To address these problems, we first explore the characteristics of the model output. We find that the null space of the matrix composed of the output vectors of model is invariant under LL-LFEA. Based on this, we propose a new null space verification method that can resist LL-LFEA attack. This method uses a new metric, the Null Space Matching Degree (NSMD). NSMD measures the degree of match between the output matrix of the suspicious model and the null space of the protected PLM. Finally, we propose NSmark, a null space based task-agnostic black-box watermarking scheme for PLMs. NSmark  uses identity information to generate all watermark-related elements, and uses WER and NSMD to verify the watermark. Spread spectrum modulation technology and an extra extractor are also introduced to enhance the watermark performance.
Our contributions are summarized as follows:
(i) We analyze the threat of LFEA on output-based watermark, and propose LL-LFEA, which can destroy the watermark embedded in the output vector without affecting the performance of downstream tasks.
(ii) We find that the null space of the matrix composed of the output vectors of model is invariant under LL-LFEA and thus propose a new null space verification method NSmark  that can resist LL-LFEA. Notably, NSmark  is task-agnostic that uses both new null space verification and signature verification to resist LL-LFEA.
(iii) We conduct comprehensive experiments by applying NSmark  to various models of both pre-training and downstream tasks. Experimental results demonstrate the effectiveness, fidelity, reliability and robustness of NSmark.
2 Related Work
Watermarking for PLMs. With the rise of pre-training in NLP, recent work has explored watermarking specific to PLMs. BadPre Jia et al. (2022) introduces a task-agnostic backdoor attack only for MLM-based PLMs. Hufu Xu et al. (2024) introduces a modality-agnostic approach for pre-trained Transformer models using the permutation equivariance property. Explanation as a Watermark Shao et al. (2024) addresses the limitations of backdoor-based techniques by embedding multi-bit watermarks into feature attributions using explainable AI. Shen et al. (2021); Zhang et al. (2023) propose task-agnostic backdoor attacks by assigning high-dimensional vectors as trigger set labels, but their effectiveness is sensitive to downstream classifier initialization. Wang and Kerschbaum (2021) introduced an auxiliary neural network for watermark embedding using weights from the main network. Wu et al. (2022) proposed a task-agnostic embedding loss function, but didn’t consider the need for triggers to reflect the model owner’s identity. Cong et al. (2022) introduced a black-box watermarking scheme for PLMs, but its applicability is limited due to the discrete nature of word tokens. Unfortunately, these schemes are vulnerable to attacks by LFEA or LL-LFEA in principle.
Watermark Removal Attacks. DNN watermarking faces various removal attempts. Common methods include fine-tuning Adi et al. (2018) and pruning Han et al. (2015). Fine-pruning Liu et al. (2018) combines these approaches for increased effectiveness. Knowledge Distillation Hinton (2015) techniques can also inadvertently remove watermarks while reducing model size. Lukas et al. (2022) propose a new attack method called Neuron Reordering to swap neurons within the same hidden layer of a DNN to disrupt embedded watermarks in the model’s parameters. Li et al. (2023a) introduce a powerful LFEA for white-box watermarks, applying linear transformations to model parameters, effectively destroying embedded watermarks while preserving the model’s original functionality. Fraud attacks include overwriting Wang and Kerschbaum (2019) and ambiguity attacks Zhu et al. (2020) also pose a great threat to watermarks.
3 Method
3.1 Threat Model
In white-box watermarking schemes, high-dimensional model parameters are often used as watermark information. For PLMs, since the output of the last layer is a high-dimensional vector, we can use a method similar to the white-box watermarking scheme to embed watermarks in the output. However, embedding identity information into the high-dimensional output vector will face the threat of LFEA-like attacks, which is proposed to destroy watermark information embedded in model parameters by linearly transforming parameters of intermediate layers. Next, we discuss the specific form of linear isomorphism attacks in this scenario.
Assuming that the attacker knows the watermark information is embedded in the output of the PLM, and hopes to destroy the watermark at a very low attack cost (without modifying the model structure or fine-tuning the model) without affecting the normal task performance of the model. As shown in Figure 2, we give an attack method that can meet this requirement and provide a proof below. Output vector $\vec{x}$ is obtained from PLM, which is also the input of downstream model. After a series of linear and nonlinear layers, the prediction result $y$ is obtained. The attacker attempts to destroy the watermark by changing the output vector of PLM, but does not affect the final prediction result of the model, that is, the attacker hopes to change $\vec{x}$ to $\varphi(x)$ and input it into the downstream network, and corresponding prediction result $y^{\prime}=y$. The sufficient condition for this result is that the modification of the PLM output vector can be offset after passing through the first linear layer of the downstream network. Let the parameter of the first linear layer of the downstream network be $W$, then the attacker hopes $W^{\prime}\varphi(\vec{x})=W\vec{x}$, then we have $\varphi(\vec{x})=W^{\prime,\dagger}W\vec{x}=Q\vec{x}$, where $Q=W^{\prime,\dagger}W$ and $W^{\prime,\dagger}$ is the pseudo-inverse of $W^{\prime}$ Li et al. (2023a). In order not to lose information in linear transformation (otherwise downstream tasks will be affected uncontrollably), $Q$ must be a reversible matrix. This proves that attacker can perform a linear transformation on $\vec{x}$ to destroy the watermark embedded in the output vector without affecting the performance of downstream tasks. We call this attack the Last-Layer Linear Functionality Equivalence Attack (LL-LFEA).
Figure 2: The schematic diagram of model inference flow before and after LL-LFEA attack.
3.2 Null Space Verification Theory
LL-LFEA performs linear transformation on output vector of PLM and can destroy the watermark embedded in it, which means previous watermark verification methods will be significantly affected. We find that the null space of the matrix composed of the output vector is invariant under the LL-LFEA attack, and propose to use the null space matching degree to verify whether the model is embedded with watermarks.
Theorem 1.
Before and after LL-LFEA, the null space of the output matrix of PLM remains unchanged for the same input set. (Proofs in Appendix A.1)
Figure 3: The overall workflow of NSmark.
Therefore, even if the watermark based on the digital string is corrupted, we can still verify model ownership using the null space of the output matrix.
3.3 Null Space Match Degree (NSMD)
We define NSMD by introducing the distribution of elements in a matrix, which is obtained by matrix multiplication of the output matrix $A$ of any PLM without watermark and the null space matrix $N$ of $f_{wm}$. In $H_{(n\times p)}=A_{(n\times m)}\times N_{(m\times p)}$, $H_{i,j}=\alpha_{i}\cdot\beta_{j}$ is the dot product of the i-th row vector of $A$ and the j-th column vector of $N$. We define NSMD of $A$ and $N$ as:
$\text{NSMD}(A,N)=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{p}\sqrt{|H_{i,j}|}.$ (1)
Furthermore, we give a detailed analysis of estimation of NSMD (in Appendix A.2). For example, if $n=768$ and $p=1500$, we have NSMD $>27.48$. If $N$ is the null space matrix of $A$, NSMD is a minimum value close to 00. This difference is amplified by the process of calculating the square root, resulting in a significant difference between whether $A$ and $N$ are matched. We use this difference to distinguish whether the model is embedded with a watermark.
3.4 Overall Framework of NSmark
As shown in Figure 3, NSmark  scheme includes three modules: watermark generation, watermark embedding, and watermark verification:
3.4.1 Watermark Generation
The workflow of watermark generation is shown in Algorithm 1. We hope that the generated watermark contains the owner’s identity information. First, the digital signature $sig=\textbf{Sign}(m)$ is generated from the identity information message $m$. In order to ensure that the trigger $t$ has a unique mapping relationship with the $sig$, only one trigger is used. We use the trigger generation algorithm $\textbf{Encode}(\cdot)$ introduced in Li et al. (2023b) to obtain $t=\textbf{Encode}(sig,n=1)$. $t$ is inserted into clean sample $x$ of dataset $D$ to form a trigger set $D_{T}$.
In order to defend against ambiguous attacks, the verification trigger set $D_{V}$ used for null space verification also needs to be generated based on $sig$. A candidate pool $D_{NS}$ for generating null space verification data sets needs to be published, and then a fixed number of samples are selected from the $D_{NS}$ based on the digital signature as the verification data set $D_{V}$. We define the verification data set selection algorithm as $\textbf{Select}(sig)\rightarrow D_{V}$, which must be a deterministic algorithm, that is, for the same input, there must be the same output. In addition, we hope that the algorithm has different outputs for different inputs. Therefore, we choose a hash function and use a one-way hash chain to generate $D_{V}$. We hope that the index repetition rate obtained by different hash value mappings is low, so we hope that the data set $D_{NS}$ is as large as possible. The specific process of the $\textbf{Select}(\cdot)$ algorithm is shown in Algorithm 2.
In order to improve the robustness of the watermark, we introduce spread spectrum modulation technology as Feng and Zhang (2020). Spread spectrum modulation technology uses redundant bits to represent the original information. Figure 4 shows an example of 3×3$\times$3 × spreading. Please refer to the Appendix A.5.1 for the specific process $\textbf{SM}(sig)\rightarrow sig_{wm}$.
Figure 4: Example diagram of spread spectrum modulation.
3.4.2 Watermark Embedding
Before the training starts, make a copy of $f_{wm}$ as the frozen reference model $f_{ref}$. Then use the clean data set $D$ and the trigger set $D_{T}$ to train $f_{wm}$ and the extractor $E$. When taking $\{D,D_{T}\}$ as input, $f_{wm}$ will output $\{V,V^{T}\}$ and $f_{ref}$ will output $\{V_{ref},V_{ref}^{T}\}$, respectively. For $V^{T}$, $E$ maps it to obtain the signature $sig_{wm}$, and for $\{V,V_{ref},V_{ref}^{T}\}$, $E$ maps them to random vectors. After the training is completed, $f_{wm}$ is embedded with watermark. Then using $D_{V}$ as input, the output vectors are concatenated into a matrix $A$, and the corresponding null space matrix $N$ of $A$ is calculated as part of the key.
Three networks are involved in watermark embedding: the model $f_{wm}$ to be embedded with watermark, the reference model $f_{ref}$, and the extractor model $E$. Compared with directly embedding $sig$ into the output vector of $f_{wm}$, adding $E$ to map can reduce the side effect of watermark on the original performance. Watermark capacity is increased at the same time. We use mean square error loss (MSE) and similarity function $\operatorname{sim}$ to implement the above training process:
$L_{\text{match}}=\frac{1}{|D_{T}|}\sum_{x\in D_{T}}\operatorname{MSE}\left(E\left(f_{wm}(x)\right),{sig}_{sm}\right),$ (2)
$L_{\text{random }}$ $=\frac{1}{|D|}\sum_{x\in D}\operatorname{sim}\left(E\left(f_{wm}(x)\right),sig_{sm}\right)^{2}$ (3)
$+\frac{1}{|D_{T}|}\sum_{x\in D_{T}}\operatorname{sim}\left(E\left(f_{{ref}}(x)\right),sig_{sm}\right)^{2}$
$+\frac{1}{|D|}\sum_{x\in D}\operatorname{sim}\left(E\left(f_{{ref}}(x)\right),sig_{sm}\right)^{2}.$
We use cosine similarity as the $\operatorname{sim}$ function. The complete loss function of $E$ is $L_{\text{Extractor}}=\lambda_{1}L_{\text{match}}+(1-\lambda_{1})L_{\text{random}}$. When training $E$ using the loss, only the parameters of $E$ are trainable. The loss of $f_{wm}$ also consists of two parts: $L_{wm}=\lambda_{2}L_{match}+(1-\lambda_{2})L_{0}$. The content of this $L_{match}$ is the same as $L_{match}$ of $E$, but only the parameters of $f_{wm}$ are updated at this time, and $L_{0}$ is the original training loss function of PLM. During training, $E$ and $f_{wm}$ are trained alternately.
3.4.3 Watermark Verification
To better defend against possible attacks, NSmark  uses two metrics together to verify the ownership: Watermark Extracting Rate (WER) and NSMD. Model owner needs to submit $key=(sig,E,N)$ to Certification Authority (CA). CA generates $t,D_{V},sm$ using $sig$. Input $D_{V}$ to the suspicious model $f_{susp}$ to get the output vector $A_{susp}$, and pass $A_{susp}$ through $E$ to get the mapped vector $O_{susp}$. Then WER is calculated after despread spectrum. Despread spectrum is the inverse process of spread spectrum (detailed process in Appendix A.5.2). At last the signature is extracted as $sig^{\prime}=\{a^{\prime}_{i}|a^{\prime}_{i}\in\{-1,0,1\},i\in[0,n-1]\}$. WER is defined as the proportion of $sig$ and $sig^{\prime}$ having the same value:
$\text{WER}=\frac{1}{n}\sum_{i=0}^{n-1}\left[a_{i}=a_{i}^{\prime}\right].$ (4)
[…]
NSMD is calculated using $A_{susp}$ and $N$ by Equation 1. We define two thresholds, and whether WER$>T_{W}$ will be firstly verified. If it fails, whether NSMD$<T_{N}$ will be further considered in case of LL-LFEA.
4 Experiments
4.1 Experimental Setup
Datasets. We use WikiText-2 Merity et al. (2017) for pre-training and watermark embedding. To evaluate the performance on downstream tasks, we select many text classification datasets: SST-2 and SST-5 Socher et al. (2013) for sentiment analysis, Lingspam Sakkis et al. (2003) for spam detection, OffensEval Zampieri et al. (2019) for offensive language identification, and AG News Zhang et al. (2015) for news classification.
Metric $f_{wm}$ $f_{clean}$
BERT RoBERTa DeBERTa XLNet BERT RoBERTa DeBERTa XLNet
WER 1.00 1.00 1.00 1.00 0.00 0.00 0.00 0.03
NSMD $2.94\times 10^{-6}$ $2.53\times 10^{-6}$ $2.91\times 10^{-6}$ $2.90\times 10^{-6}$ 60.95 61.24 87.88 76.74
Table 1: Effectiveness of NSmark  on different PLMs.
Models. For PLMs, we use the base versions of BERT Kenton and Toutanova (2019), RoBERTa Liu (2019), DeBERTa He et al. (2020) and XLNet Yang (2019). All pre-trained weights are from HuggingFace [^1] https://huggingface.co/. The extractor network is a three-layer linear network with hidden layers of $2048$ and $1024$ neurons. The input dimension matches the output dimension of the PLM. The output dimension matches the size of $sig_{sm}$.
Watermark settings and training details. We select a string containing owner information as the message $m$, e.g., "BERT is proposed by Google in 2018". The length of $sig$ is $256$ and is spread-spectrum with a spreading factor $k=3$, resulting in a $768$-bit $sig_{sm}$. We use the labeled downstream dataset SST-2 as the candidate pool $D_{NS}$. $q$, the length of $D_{V}$, is $1500$. The trigger is inserted to random positions for $5$ times in trigger set. When performing watermark embedding, $\lambda_{1}=0.5$ and $\lambda_{2}=0.2$ in $L_{Extractor}$ and $L_{wm}$. Batchsize is $4$, and learning rates for both $f_{wm}$ and $E$ are $10^{-4}$. $f_{wm}$ and $E$ are trained alternately for $10$ epochs. When fine-tuning on downstream tasks, the learning rate is $2\times 10^{-5}$ and batchsize is $8$ for $3$ epochs.
Metrics. As mentioned before, We define two metrics to verify the model’s identity: WER and NSMD. Besides, we adopt accuracy (ACC, in %) to measure the performance of PLM on downstream tasks.
4.2 Performance Evaluation
4.2.1 Effectiveness
Effectiveness means that the watermark can achieve expected effect in verification. Ideally, $sig^{\prime}$ extracted from the watermarked model should be consistent with the original $sig$, and the output matrix of $f_{wm}$ for $D_{V}$ should completely match $N$ stored in the $key$, which means WER = $1$ and NSMD = 00. Table 1 shows the results of $f_{wm}$ embedded with watermark and $f_{clean}$ without watermark. It can be seen that for different watermarked PLMs, WER is $1$, and NSMD is close to 00. This proves the effectiveness of NSmark. Comparing the values of $f_{wm}$ and $f_{clean}$ shows that WER and NSMD will change obviously after the watermark is embedded. Although NSMD of different PLMs is different in value, they are all far from 00. Through these results, we can preliminarily define verification thresholds of WER and NSMD as $T_{W}=0.6$ and $T_{N}=43$, which are 0.6×0.6$\times$0.6 × the average gaps. Thresholds can be further adjusted according to different models and task types.
4.2.2 Reliability
The watermark key is a triple $key=(sig,E,N)$. Next, we analyze whether the watermark can be successfully verified if an attacker provides an incorrect $key$.
Figure 5: The impact of the correctness of trigger $t$ and signature $sig$ on WER and NSMD.
Wrong signature $sig$. The trigger $t$ and the output of $f_{wm}$ are related to $sig$, but as $sig$ and $t$ are not a one-to-one mapping relationship, there are situations where only one of $sig$ and $t$ is correct. Figure 5 shows all possible scenarios.
For $f_{wm}$, (1) when trigger is wrong ($t_{w}$) and signature is correct ($sig_{c}$), WER = 0, NSMD $>T_{N}$. This means that $f_{wm}$ has learned the relationship between $sig$ and $t$. Whether $t$ is correct determines whether $f_{wm}$ can produce the expected output, which in turn affects both the calculation of WER and NSMD. (2) When trigger is correct ($t_{c}$) and signature is wrong (considering the most dangerous scenario $sig_{w}$ only consists of $\{-1,1\}$), WER $\approx 0.5$. This is because $t_{c}$ leads to the right $sig^{\prime}$, and its expectation of WER with a random $\{-1,1\}$ string is $0.5$. As the output matrix is correctly generated by $f_{wm}$ based on $t_{c}$, NSMD = 00 in this case. (3) When both trigger and signature are wrong ($t_{w}$ and $sig_{w}$), WER = 00, NSMD > $T_{N}$, indicating that the watermark cannot be correctly verified without providing the correct key. (4) For model without embedded watermarks $f_{clean}$, watermarks cannot be extracted even if the correct key is provided.
Setting BERT RoBERTa DeBERTa XLNet
$f_{wm}+E_{c}$ 1.00 1.00 1.00 1.00
$f_{wm}+E_{w}$ 0.00 0.00 0.00 0.13
Table 2: WER of different extractor $E$.
Wrong extractor $E$. Since $E$ is not involved in the calculation of NSMD, we only analyze the impact of $E$ on WER. As shown in Table 2, when $E$ is wrong ($E_{w}$), WER is close to 00, indicating that wrong $E$ is unable to extract the watermark.
Setting BERT RoBERTa DeBERTa XLNet
$f_{wm}+N_{c}$ $2.94\times 10^{-6}$ $2.53\times 10^{-6}$ $2.91\times 10^{-6}$ $2.90\times 10^{-6}$
$f_{wm}+N_{r}$ 3167.81 3171.58 3182.79 3117.61
$f_{wm}+N_{s}$ 1001.75 1002.94 1006.49 985.87
Table 3: NSMD of different null space matrix $N$.
Metric BERT RoBERTa DeBERTa XLNet
$f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$
WER 1.00 0.27 1.00 0.07 1.00 0.27 1.00 0.00
NSMD $2.94\times 10^{-6}$ 0.06 $2.53\times 10^{-6}$ 0.04 $2.91\times 10^{-6}$ 0.07 $2.90\times 10^{-6}$ 0.05
Table 4: Impact of LL-LFEA on watermark performance.
Wrong null space $N$. Since $N$ is not involved in the calculation of WER, we only analyze its impact on NSMD. As shown in Table 3, $N_{r}$ is a randomly generated matrix with the same dimension as $N$ and each element is distributed between $[0,1]$. It can be seen that NSMD is very large at this time. However, if the attacker knows the watermark algorithm, a $N_{s}$ with extremely small elements can be generated. In this case NSMD might meet the verification requirements. This indicates that NSMD cannot be used independently to verify watermark.
4.2.3 Fidelity
Metric Model Setting SST-2 SST-5 Offenseval Lingspam AGnews
ACC BERT $F_{wm}$ 91.40 52.62 85.12 99.14 93.95
$F_{clean}$ 91.63 53.03 84.07 99.66 94.38
RoBERTa $F_{wm}$ 92.55 54.71 84.30 99.66 94.43
$F_{clean}$ 94.04 56.15 84.88 100.00 94.72
DeBERTa $F_{wm}$ 93.00 55.48 83.02 99.31 94.61
$F_{clean}$ 93.58 57.65 85.12 99.31 94.84
XLNet $F_{wm}$ 88.65 42.67 81.98 99.14 93.29
$F_{clean}$ 93.58 53.62 84.65 99.31 94.07
WER BERT $F_{wm}$ 1.00 1.00 1.00 0.94 1.00
$F_{clean}$ 0.00 0.00 0.00 0.00 0.00
RoBERTa $F_{wm}$ 0.98 1.00 1.00 0.72 0.99
$F_{clean}$ 0.00 0.00 0.00 0.00 0.00
DeBERTa $F_{wm}$ 1.00 1.00 1.00 0.80 0.88
$F_{clean}$ 0.00 0.00 0.00 0.00 0.00
XLNet $F_{wm}$ 1.00 1.00 1.00 0.92 1.00
$F_{clean}$ 0.00 0.00 0.00 0.01 0.00
NSMD BERT $F_{wm}$ 29.77 25.29 22.52 21.96 24.37
$F_{clean}$ 72.97 70.06 66.65 69.90 61.59
RoBERTa $F_{wm}$ 50.17 30.97 25.15 26.78 28.43
$F_{clean}$ 74.75 74.48 69.90 65.06 75.54
DeBERTa $F_{wm}$ 31.89 25.74 23.52 27.23 37.68
$F_{clean}$ 80.81 76.72 72.41 68.49 76.33
XLNet $F_{wm}$ 24.12 23.52 25.29 24.06 26.20
$F_{clean}$ 76.30 74.75 69.84 78.09 74.97
Table 5: The performance of models w/o watermark fine-tuned on downstream tasks.
We hope that NSmark  does not affect the performance on original tasks. Thus we add a downstream network to $f_{wm}$, and fine-tune the whole model $F_{wm}$ with downstream dataset. $F_{clean}$ without watermark is fine-tuned as baseline. Table 5 shows that watermark has almost no impact on the performance of the model on the original task.
4.2.4 Defense against LL-LFEA
Defense against LL-LFEA. When designing NSmark, we focus on resisting LL-LFEA, and propose null space verification using NSMD. Table 4 shows the impact of the LL-LFEA on the watermark verification, where $f_{LL-LFEA}$ denotes the model $f_{wm}$ attacked by LL-LFEA. Experiments show that after LL-LFEA, WER drops significantly as we discussed in §$\S$ 3.1, but NSMD is still close to 00, proving that NSMD is an effective indicator for LL-LFEA. Furthermore, after applying LL-LFEA, the attacker may add a network to $f_{LL-LFEA}$ and fine-tune it for downstream tasks. We present results in Appendix B.1.
Recovery of WER. In LFEA Li et al. (2023a), a method is proposed to recover the watermark. We revise this method to recover $f_{rec}$ from $f_{LL-LFEA}$. Specifically, assume that the output matrix of $f_{wm}$ is $A_{1(n\times m)}$ as Proof A.1. After attacked with $Q_{(n\times n)}$, the output matrix turns to $A_{2}=Q\times A_{1}$. Therefore, an estimate of $Q$ can be obtained as $Q^{\prime}=A_{2}\times A_{1}^{-1}$. If $m\neq n$, then $A_{1}$ is not reversible and $Q^{\prime}=A_{2}\times A_{1}^{T}\times(A_{1}\times A_{1}^{T})^{-1}$. Then we perform an anti-attack transformation on $f_{LL-LFEA}$, that is, multiply all the outputs of $f_{LL-LFEA}$ by $Q^{\prime}$ to get $f_{rec}$. Figure 6 shows that after recovery, WER is significantly improved, indicating that such linear attacks are recoverable. In all cases, NSMD is quite small, proving that NSMD is invariant to LL-LFEA. In the recovery algorithm in Li et al. (2023a), both the attacker and the model owner might use such an algorithm to claim that they are the owner of the model, which will cause verification ambiguity. However, our proposed NSMD is invariant under LL-LFEA, thus as long as the timestamp information is added to $key$ tuple, the ownership can be reliably verified according to the time sequence of the model and $key$ release.
Figure 6: Changes of WER and NSMD before and after LL-LFEA attack and recovery.
4.2.5 Robustness
The robustness of watermark refers to whether the watermark can be effectively verified after watermark removal attacks. Next, we will analyze the robustness of NSmark  against fine-tuning, pruning, fine-pruning and overwriting attacks.
Robustness against fine-tuning. Table 5 shows the WER and NSMD results after fine-tuning on downstream tasks. $F_{wm}$ and $F_{clean}$ are obtained same as §$\S$ 4.2.3. In most cases, WER is still very high, indicating that the embedded $sig$ can still be effectively extracted after downstream fine-tuning. However, WER of RoBERTa on Lingspam task is relatively low. This is because samples in this dataset are too long to process, and the output vectors after fine-tuning have large changes, which together lower the WER. Nevertheless, compared with $F_{clean}$, there is still obvious discrimination. Therefore, for complex tasks, the verification threshold $T_{W}$ can be slightly lowered.
Figure 7: Impact of pruning attacks on watermark performance.
Robustness against pruning and fine-pruning. Pruning is a commonly used model compression method and is often used to destroy the watermark embedded in the model. Referring to Han et al. (2015); Shao et al. (2024), we sort the parameters of each layer in PLM, then set different fractions of parameters with the smaller absolute value to 00. Figure 7 shows that when the pruning rate is less than or equal to $0.8$, WER is very high. When the pruning rate is less than or equal to $0.6$, NSMD does not change significantly, and even when the pruning rate is as high as $0.9$, NSMD is still distinguishable. This shows that the embedded watermark is robust to pruning attack.
Usually, pruning will affect the performance of the model on the original task, and the original task accuracy will be restored through fine-tuning (fine-pruning). We present corresponding results in Appendix B.2.
Model Downstream dataset ACC WER NSMD
$f_{ow}$ 1.00 22.76
$F_{ow}$ SST-2 92.32 0.98 48.77
SST-5 50.54 1.00 36.47
Offenseval 84.77 1.00 36.61
Lingspam 99.31 0.62 28.87
AGnews 93.51 1.00 32.49
Table 6: Impact of overwrite attacks on watermark performance.
Robustness against overwriting. Overwriting means attacker embeds his own watermark into a model that has already been watermarked in the same way. This may destroy the original watermark. We simulate this process to obtain $f_{ow}$, then add a downstream network and fine-tune to obtain $F_{ow}$. We test the original watermark as shown in Table 6. Overwriting attack has little effect on ACC and WER except on Lingspam. Meanwhile, it has a similar impact on NSMD as fine-tuning.
5 Further Analysis
Model Metric $f_{wm}$ $F_{wm}$
SST-2 SST-5 Offenseval Lingspam AGnews
BERT NSMD $2.94\times 10^{-6}$ 29.77 25.29 22.52 21.96 24.37
$\text{NSMD}_{c}$ $3.01\times 10^{-6}$ 76.20 73.39 64.55 66.60 74.75
RoBERTa NSMD $2.53\times 10^{-6}$ 50.17 30.97 25.15 26.78 28.43
$\text{NSMD}_{c}$ $2.54\times 10^{-6}$ 74.37 71.75 64.50 65.90 74.95
DeBERTa NSMD $2.91\times 10^{-6}$ 31.89 25.74 23.52 27.23 37.69
$\text{NSMD}_{c}$ $2.98\times 10^{-6}$ 74.31 71.83 64.48 50.73 73.67
XLNet NSMD $2.90\times 10^{-6}$ 24.12 23.52 25.29 24.06 26.20
$\text{NSMD}_{c}$ $3.00\times 10^{-6}$ 68.86 55.43 41.23 24.88 38.02
Table 7: Impact of using a validation set without triggers on watermarking performance.
Next we discuss the necessity of using the verification set with trigger to calculate NSMD. In §$\S$ 3.4.1 of watermark generation, we introduce the selection and generation method of the verification trigger set. In order to explain why uses the trigger set instead of the clean set to form $D_{V}$, we compare the effects of trigger set (NSMD) with clean set ($\text{NSMD}_{c}$) in Table 7. It shows that for $f_{wm}$, NSMD is always close to 00, but the model shows a huge difference after downstream fine-tuning. In this case, $\text{NSMD}_{c}$ is significantly higher than NSMD, indicating that the output matrix no longer matches the original null space matrix. This is because only the output of triggered set is mapped to $sig$ through $E$. Results in Table 5 show that fine-tuning has little effect on WER, indicating that fine-tuning has little effect on the output representation of trigger set. This means the output matrix of trigger set changes little, so it still matches original $N$, and NSMD is always close to 00. However, after fine-tuning, the output representation corresponding to the clean data set will change significantly to achieve better performance on different downstream tasks, which causes the output matrix no longer match the original $N$, causing $\text{NSMD}_{c}$ to significantly increase. Therefore, the trigger set is needed for verifying null space.
6 Conclusion
This paper proposes NSmark, a black-box watermark framework for ownership verification using the output of PLM. We first analyze and introduce LL-LFEA, and propose a solution that can use null space invariance for watermark verification. We conduct an overall design from three aspects: watermark generation, watermark embedding, and watermark verification. Two indicators, WER and NSMD, are used to jointly verify the existence and identity of the watermark. Experiments prove the effectiveness, reliability and fidelity of NSmark, and it has satisfactory performance under various attacks. With the cooperation of two verification methods, it works a secure and robust watermarking scheme.
7 Limitations
Our approach has limitations in two main aspects. First, our method only focuses on LMs as we use the vocab to build connection between $sig$ and trigger. Solutions applicable to other models (such as visual or multi-modal models) need further exploration. Second, further research is needed on this basis to further protect the downstream network.
Appendix A Additional Details of Theory and Algorithm
A.1 Proofs of Null Space Verification Theory
Proof.
The null space $N(A)$ of matrix $A_{(a\times b)}$ is the set of all $b$-dimensional vectors $x$ that satisfy $Ax=\vec{0}$ Axler (2015). That is, $N(A)=\{x\in\mathbb{R}^{b},Ax=\vec{0}\}$. Using $f_{wm}$ to denote PLM embedded with watermark, assuming $A_{1(n\times m)}=\{f_{wm}(x),x\in D_{T}\}$ is the matrix concatenated from the output vectors, where $D_{T}$ is the verification dataset with watermark trigger, $m$ is the size of $D_{T}$ and $n$ is the dimension of the output vector of the last layer of PLM. Let the null space matrix of $A_{1}$ be $N_{1}$, then $A_{1}\times N_{1}=\textbf{0}$.
After performing LL-LFEA, assuming the new output matrix of $f_{wm}(D_{T})$ is $A_{2(n\times m)}$, according to Sec.3.1, we have $A_{2}=Q\times A_{1}$. Then $A_{2}\times N_{1}=(Q\times A_{1})\times N_{1}=Q\times(A_{1}\times N_{1})=\textbf{0}$, which means $N_{1}$ belongs to the null space matrix of $A_{2}$. As $Q$ is a reversible matrix, then $rank(A_{1})=rank(A_{2})$, and the null space of $A_{1}$ and $A_{2}$ have the same dimension. It can be concluded that $N_{1}$ is also the null space matrix of $A_{2}$. ∎
A.2 Estimation of NSMD
We define NSMD by introducing the distribution of elements in a matrix, which is obtained by matrix multiplication of the output matrix $A$ of any PLM without watermark and the null space matrix $N$ of $f_{wm}$. In $H_{(n\times p)}=A_{(n\times m)}\times N_{(m\times p)}$, $H_{j,j}=\alpha_{i}\cdot\beta_{j}$ is the dot product of the i-th row vector of $A$ and the j-th column vector of $N$. It is proposed that the approximate distribution of the angle between $n$ random uniformly distributed unit vectors in space $\mathbb{R}^{m}$ Cai et al. (2013). In space $\mathbb{R}^{m}$, given two random vectors uniformly distributed on the unit sphere, the angle $\theta$ between the two random vectors converges to a distribution whose probability density function is:
$f(\theta)=\frac{1}{\sqrt{\pi}}\cdot\frac{\Gamma(\frac{m}{2})}{\Gamma(\frac{m-1}{2})}\cdot(\sin\theta)^{m-2},\theta\in[0,\pi].$ (5)
When $m=2$, $f(\theta)$ is uniformly distributed on $[0,\pi]$; when $m>2$, $f(\theta)$ has a single peak at $\theta=\frac{\pi}{2}$. When $m>5$, the distribution of $f(\theta)$ is very close to the normal distribution. Most of the $C_{m}^{2}$ angles formed by $m$ randomly uniformly distributed unit vectors are concentrated around $\frac{\pi}{2}$, and this clustering will enhance with the increase of the dimension $m$, because if $\theta\neq\frac{\pi}{2}$, then $(\sin\theta)^{m-2}$ will converge to 00 faster. This shows that in high-dimensional space, two randomly selected vectors are almost orthogonal.
We further derive the distribution of the dot product of two random vectors uniformly distributed and independently selected on the unit ball in space $\mathbb{R}^{m}$. Let $\alpha$ and $\beta$ be unit vectors and $\theta$ be the angle between them, then $\alpha\cdot\beta=\cos(\theta)$. It is known that $\theta$ obeys the probability distribution $f(\theta)$, then the probability density function of $y=\alpha\cdot\beta=\cos(\theta),y\in[-1,1]$ is:
$\begin{split}g(y)=&g(\cos(\theta))=f(\arccos(\cos(\theta)))\cdot\\ &|d(\arccos(\cos(\theta)))/d(\cos(\theta))|,\end{split}$ (6)
where $d(\arccos(\cos(\theta)))/d(\cos(\theta))=-1/\sqrt{(1-\cos^{2}(\theta))}$ is the derivative of the inverse cosine function. It can be inferred that:
$g(y)=g(\cos(\theta))=f(\theta)/\sqrt{(1-\cos^{2}(\theta))}.$ (7)
Further, we analyze the mathematical expectation and variance of $Y=\cos(\Theta)$. The mean is:
[…]
Note $k_{m}=\frac{1}{\sqrt{\pi}}\cdot\frac{\Gamma\left(\frac{m}{2}\right)}{\Gamma\left(\frac{m-1}{2}\right)}$, then:
[…]
Its variance is:
[…]
$=\frac{2}{\sqrt{\pi}}\cdot\frac{\Gamma\left(\frac{m}{2}\right)}{\Gamma\left(\frac{m-1}{2}\right)}\cdot\left(I_{m-2}-I_{m}\right),$
where:
[…]
$m$ 10 20 300 768 1024 100000
DY 0.15667 0.11217 0.029302 0.018323 0.015870 $7.1830\times 10^{-6}$
Table 8: The value of DY in different dimensions $m$.
Figure 8: The relationship between the variance DY and the spatial dimension $m$.
As $m$ increases, $DY$ gradually approaches 00. Figure 8 shows the relationship between the DY and the $m$, and Table 8 shows the specific values of the variance when $m$ takes specific values. Under the common output dimension of PLM, that is, when $m=1000$ or so, $DY$ is still a distance to 00.
Because the multiplication of the output matrix $A_{1}$ of the model embedded with watermark and its null space $N_{1}$ is exactly 00, while the variance of the elements obtained by the multiplication of the output matrix of other irrelevant models and $N_{1}$ is different from 00, we use and amplify this gap to define a new verification indicator - Null Space Match Degree (NSMD) for watermark verification.
For an output matrix $A_{(n\times m)}$ and a null space matrix $N_{(m\times p)}$, we first normalize all row vectors $\alpha_{i},i\in[1,n]$ of $A$ and all column vectors $\beta_{j},j\in[1,p]$ of $N$ so that $\alpha$ and $\beta$ are distributed on the unit sphere, and then calculate the $H_{n\times p}=A\times N$. We define NSMD of $A$ and $N$:
$\text{NSMD}(A,N)=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{p}\sqrt{|H_{i,j}|}.$ (12)
As $\sqrt{|h_{i,j}|}\in[0,1]$ and $DY=0$, we have
$\text{NSMD}(A,N)$ $>\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{p}H_{i,j}^{2}$ (13)
$=p\cdot EY^{2}$
$=p\cdot(DY+(EY)^{2})$
$=p\cdot DY.$
Furthermore, $\text{NSMD}(A,N)>p\cdot DY$. For example, if $n=768$ and $p=1500$, we have NSMD $>27.48$.
A.3 Trigger Generation Algorithm
Algorithm 1 Trigger Generation Algorithm
Input: owner’s private key $K_{pri}$, identity information message $m$
Output: digital signature $sig$, trigger word $t$, verification set $D_{V}$
1:  $sig\leftarrow\textbf{Sign}(m,K_{pri})$.
2:  $t\leftarrow\textbf{Encode}(sig,n=1)$
3:  $sig_{sm}\leftarrow\textbf{SM}(sig)$
4:  $D_{V}\leftarrow\textbf{Select}(sig)$
5:  return $sig,sig_{sm},t,D_{V}$
The trigger generation algorithm is shown in Algorithm 1.
A.4 Select Algorithm
Algorithm 2 Select Algorithm
Input: digital signature $sig$, $|D_{V}|=q$, candidate data pool $D_{NS}$
Output: verification set $D_{V}$
1:  initialize $D_{V}\leftarrow[\ ]$.
2:  $h_{0}\leftarrow\textbf{Hash}(sig)$
3:  for $i=1$ to $q$ do
4:     $h_{i}\leftarrow\textbf{Hash}(h_{i-1})$
5:     $idx_{i}\leftarrow h_{i}\%\ len(D_{NS})$
6:     $D_{V}.\text{append}(D_{NS}[idx_{i}])$
7:  end for
8:  return $D_{V}$
The $\textbf{Select}(\cdot)$ algorithm is shown in Algorithm 2.
A.5 Process of Spread Spectrum Modulation and Despread Spectrum
A.5.1 Spread Spectrum Modulation
Assume that the digital signature $sig$ is $n$ bits, $sig=\{a_{i}|a_{i}\in\{-1,1\},i\in[0,n-1]\}$, and set the spreading factor to $k$. Expand $sig$ horizontally by $k$ times to obtain $sig_{repeat}=\{ra_{j}|ra_{j}=a_{i},i=j\ mod\ n,ra_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$. Input $sig$ as a seed into the pseudo-random generator to obtain the key $sm=\{b_{j}|b_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$ for spread spectrum modulation. Use $sm$ to modulate $sig_{repeat}$ to obtain the spread spectrum modulated digital signature $sig_{sm}=\{sa_{j}|sa_{j}=ra_{j}\times b_{j},j\in[0,k\times n-1]\}$. Figure 4 shows an example of 3×3$\times$3 × spreading.
A.5.2 Despread Spectrum
Let the mapping vector output by $E$ be $O=\{o_{j},j\in[0,k\times n-1]\}$, modulate it with $sm=\{b_{j}|b_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$ to get $O_{repeat}=\{ro_{j}|ro_{j}=o_{j}/b_{j},j\in[0,k\times n-1]\}$, then quantify $O_{repeat}$ to get $O_{quan}=\{qo_{j}|qo_{j}\in\{-1,0,1\},j\in[0,k\times n-1]\}$. At last the signature is extracted by counting the number of $n$ positions that appears most often in $k$ copies. $sig^{\prime}=\{a^{\prime}_{i}|a^{\prime}_{i}\in\{-1,0,1\},i\in[0,n-1]\}$. The quantification method is shown as follows:
$qo_{j}=\begin{cases}1&,0.5<ro_{j}<1.5\\ -1&,-1.5<ro_{j}<-0.5.\\ 0&,\text{otherwise}\end{cases}$ (14)
Appendix B Additional Experimental Results and Analyses
B.1 Defense against LL-LFEA+finetuning
Metric Model SST-2 SST-5 Offenseval Lingspam AGnews
WER BERT 0.29 0.29 0.28 0.31 0.37
RoBERTa 0.47 0.07 0.07 0.35 0.35
DeBERTa 0.30 0.28 0.29 0.29 0.42
XLNet 0.00 0.00 0.00 0.01 0.01
NSMD BERT 15.43 15.06 12.95 12.56 13.17
RoBERTa 47.39 17.96 14.64 12.73 28.89
DeBERTa 26.86 17.51 20.38 20.77 38.38
XLNet 13.20 12.08 14.12 12.98 14.37
ACC BERT 91.17 52.22 86.04 99.48 93.80
RoBERTa 93.00 52.71 84.88 99.14 94.37
DeBERTa 94.04 51.95 82.91 99.48 93.70
XLNet 90.14 42.40 81.40 99.48 93.03
Table 9: Impact of LL-LFEA+ fine-tuning attack on watermark performance.
After applying LL-LFEA, the attacker may add a network to $f_{LL-LFEA}$ and fine-tune it for downstream tasks. We hope the model after the LL-LFEA+fine-tuning attack can still maintain the watermark. Table 9 shows results on different PLMs and different downstream tasks are not exactly the same. WER of different models has decreased significantly to varying degrees. Most NSMDs are still below the threshold, but RoBERTa and DeBERTa change more on SST-2, which is generally consistent with that of fine-tuning without LL-LFEA attack (Table 5). Through ACC, we can find that LL-LFEA attack does not affect the performance of the model on the original task.
B.2 Robustness against Fine-pruning
Pruning Rate 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 $f_{clean}$
ACC 52.62 52.35 51.11 51.99 52.35 51.63 52.94 52.71 50.81 51.44 53.03
WER 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.67 0.00
NSMD 25.29 20.13 18.45 20.76 21.18 21.19 25.37 28.78 42.07 50.81 70.06
Table 10: Impact of fine-pruning attack on watermark performance.
Figure 9: The t-SNE visualization of output feature vectors of watermarked models. (1) Left column: $f_{wm}$ on WikiText; (2) Middle column: $F_{wm}$ on SST-2; (3) Right column: $F_{wm}$ on SST-5.
Table 10 uses SST-5 as the fine-tuning dataset to show the watermark extraction effect after fine-pruning. It can be seen that the results are generally the same as Figure 7.
B.3 Feature Visualization
To demonstrate the effectiveness of our scheme, we use t-SNE to visualize the feature distribution of the watermarking model. As shown in Figure 9, the input with trigger and the input without trigger can be well separated in the output of PLM and $E$, whether for $f_{wm}$ or fine-tuned $F_{wm}$.